I owed $1,975 to a pension program in July, 1983. If interest on that amount is calculated at a rate of 8.5% annually.... how much would I owe July, 2009?
Answer by rfer(16322)

I assume it is 8.5% per yr on the unpaid balance.
A=P(1+r)^t
A=1975(1+.085)^16
A=1975(3.689)
A=$7285.22